Alice cuts a circular piece of paper into sectors of equal sizes. She rearranges all sectors in a plane without gaps or overlaps with alternating curved parts. Which of these relates the area of the new arrangement to the area of the original whole piece of paper?
1
Expert's answer
2020-04-23T18:51:42-0400

The sectors form a parallelogram with a height of "r" and base of "\pi r" . So the area of original circular piece of paper was "\pi r\times r = \pi r^2"